# Opinion formation in Wikipedia Ising networks

### v2

@@ I Introduction
The process of opinion formation in human society gains higher and higher importance with the development of social networks which start to produce an important impact on political views and elections (see e.g. [1, 2]). The statistical properties of such social networks typically have a scale-free structure as reviewed in [3, 4]. Various voter models have been proposed and studied by different groups with a development of physical concepts and their applications to sociophysics [3, 5, 6, 7, 8, 9, 10, 11, 12, 13].
Recently, we proposed the Ising Network Opinion Formation (INOF) model and analyzed its applications to Wikipedia networks for 6 language editions of 2017 [14]. This model allows to determine opinion polarization for all Wikipedia articles (or nodes) induced by two groups of nodes with fixed opposite opinions (red or blue, spin up or down). In this INOF model the initial two groups of one or a few nodes have fixed opposite opinions represented by spin up (red color) or spin down (blue color). All other nodes have initially an undecided opinion (spin zero or white color). The formation of a steady-state opinion of each node emerges as a result of an asynchronous Monte Carlo process in which an opinion of a given node $i$ is determined by a majority vote of his friends presented by spins up or down or zero from all network nodes $j$ that have links to node $i$. Such spin flips, induced by local majority votes, are done for all nodes without repetitions in random order over all $N$ nodes. These procedure is repeated up to convergence to a steady-state for a sufficiently long time $\tau$ and corresponds to a particular random path pathway realization for the order of spins to be flipped. Finally, an average over a high number of pathways realizations is done to obtain averages and distributions of the opinions for nodes or the whole network (see next section for more technical details). A somewhat similar procedure is used in the studies of problems of associative memory (see e.g. [15, 16]) even if there are significant differences from the INOF model due to the absence of certain fixed nodes and other initially white nodes and the use of positive/negative transition elements between nodes while all of them are positive for the INOF case considered here.
@@
In this work, we extend the studies of the INOF model [14] to more recent Wikipedia networks collected either in October 2024, as described in [17], or collected at 20 March 2025 and to new specific initial groups with fixed opinions. For example, we analyze the competition between Apple Inc. vs. Microsoft, Donald Trump vs. Vladimir Putin and others. Furthermore, the competition of three entries is also considered for several cases and finally, we also examine the effects of fluctuations appearing as a result of a certain effective temperature in the voting process.
@@ II Model description and data sets
[…]
Wikipedia articles correspond to network nodes and citation from a given article $j$ to another article $i$ correspond to a directed link with the adjacency matrix element $A_{ij}=1$ (and $A_{ij}=0$ in absence of a link from $j$ to $i$); multiple citations from $j$ to $i$ are considered as only one link. Then the matrix of Markov transitions is defined by $S_{ij}=A_{ij}/k_{j}$ where $k_{j}=\sum_{i}A_{ij}$ is a number of out-going links from node $j$ to any other node $i$ (such that $A_{ij}\neq 0$); for the case of dangling nodes without out-going links (i.e. with $k_{j}=0$), we simply define $S_{ij}=1/N$ implying the usual column sum normalization $\sum_{i}S_{ij}=1$ for all $j$. For later use, we also introduce the modified matrix $\tilde{S}_{ij}$ which is identical to $S_{ij}$ for $k_{j}>0$ and with $\tilde{S}_{ij}=0$ for dangling nodes.
Usually, in other typical types of network studies (see e.g. [17]) one introduces the Google matrix of the network defined as $G_{ij}=\alpha S_{ij}+(1-\alpha)/N$ where alpha is the damping factor with the standard value $\alpha=0.85$ [24, 25, 26]. Here the network nodes can be characterized by the PageRank vector which is the eigenvector of the Google matrix $G$ [24, 25, 26] with the highest eigenvalue $\lambda=1$, i.e. $GP=\lambda P=P$, and the damping factor $\alpha<1$ ensures that this vector is unique and can be computed efficiently. Its components $P(i)$ are positive and normalized to unity (${\sum_{i=1}}^{N}P(i)=1$). The network nodes $i$ can be ordered by monotonically decreasing probabilities $P(i)$ which provides the PageRank index $K$ with with highest probability at $K=1$ and smallest at $K=N$. Some results for the PageRank vector and its index for recent Wikipedia editions of 2024 can be found at [17]. However, in this work, we do not use the Google matrix neither the PageRank and focus mostly on the modified matrix $\tilde{S}$ and also the adjacency matrix $A_{ij}$ to define an asynchronous Monte Carlo process.
As in [14] a few selected nodes (wiki-articles) have assigned fixed spin values $\sigma_{l}=-1$ blue e.g. for Microsoft and $\sigma_{k}=1$ red for Apple Inc.. These specific spin nodes always keep their polarization. All other nodes $i$ are initially assigned with a white color (or spin $\sigma_{i}=0$) and have no definite initial opinion. However, once they acquire a different color red or blue (spin value $\sigma_{i}=\pm 1$) during the asynchronous Monte Carlo process they can flip only between ±1$\pm$ 1± 1 and cannot change back to the white opinion.
@@
In Eq. (1) $\sigma_{j}=1$ if the spin of node $j$ is oriented up (red color), or $\sigma_{j}=-1$ if it is oriented down (blue color), or $\sigma_{j}=0$ if the node $j$ has no opinion (if it has still its initial white value). After the computation of $Z_{i}$, the spin $\sigma_{i}$ of node $i$, is updated: it becomes $\sigma_{i}=1$ if $Z_{i}>0$, $\sigma_{i}=-1$ if $Z_{i}<0$, and remains unchanged if $Z_{i}=0$. This operation is repeated for all non-fixed nodes $i^{\prime}$ following a predetermined random order (shuffle) such that there is no repetition at this level and each spin is updated only once. Note that due to the possibility of $Z_{i}=0$ it is possible that a node $i$ keeps his initial white value $\sigma_{i}=0$. After the update the modified value of $\sigma_{i}$ is used for the computation of $Z_{i^{\prime}}$ of subsequent values $\sigma_{i^{\prime}}$.
@@
The voting process for the case of a competition between three groups of entries is an extension of this procedure and its details are be explained later.
@@ III.1 Comparison of OPA and OPS
[…]
@@
[…]
@@
[…]
@@
We also consider the INOF competition between Emmanuel Macron ($\mu=1$) vs. Marine Le Pen ($\mu=-1$) with $\mu_{0}=-0.028$ from FR Wiki2024. These corresponds to rather close vote fractions with $f_{r}=(1+\mu_{0})/2=0.4860$ for Macron and $f_{b}=(1-\mu_{0})/2=0.5140$ for Le Pen (number of pathway realizations $N_{r}=10^{6}$). The polarization opinions of 14 French political figures and top 10 French richest persons from the Forbes list 2015-2024 are presented in Appendix Table A1. The results for these 24 persons show that only two of them have a polarization in favor of Emmanuel Macron and other 22 in favor of Marine Le Pen. At the same time the values of $\Delta\mu$ are mainly located at relatively small values $|\Delta\mu|\sim 0.01$ being approximately by a factor 10 smaller of values of for the contest of Trump-Putin shown in Table 1. Due to these smaller $\Delta\mu$ values we with also compute an additional data set with $N_{r}=10^{6}$ to reduce the statistical errors (note that the FR Wiki2024 network is roughly 2 times smaller in number of nodes and links as compared to those of EN Wiki2025 thus reducing the numerical effort). Table A1 shows two columns of $\Delta\mu$ computed for $N_{r}=10^{5}$ and $N_{r}=10^{6}$ (the smaller data set is statistically independent and not included in the larger date set). The values are rather close and the small differences indicate the size of the typical fluctuations at $N_{r}=10^{5}$.
The theoretical error of the $\mu_{i}$ values is $(1-\mu_{i}^{2})/\sqrt{N_{r}}\approx 0.003$ (for $N=10^{5}$) or 0.001$\approx$ 0.001≈ 0.001 (for $N=10^{6}$). However, the more precise error estimation for $\Delta\mu_{i}=\mu_{i}-\mu_{0}$ using a subdivision of the data in 100 samples (described above after the discussion for Figure 2) provides typical statistical errors of $\Delta\mu_{i}$ for the entries in Table A1 which are roughly 3 times smaller than the theoretical error, i.e. 0.001$\approx$ 0.001≈ 0.001 (for $N=10^{5}$) or 0.0003$\approx$ 0.0003≈ 0.0003 (for $N_{r}=10^{6}$) which is due to rather strong correlations between $\mu_{i}$ and $\mu_{0}$. Therefore, despite the typical small values of $\Delta\mu_{i}\approx 0.01$ in Table A1 their relative error is mostly only 3% (for the $N_{r}=10^{6}$ data). Note, that the error 0.00230.00230.0023 of $\mu_{0}$ itself is closer to the theoretical error 0.0030.0030.003 which is confirmed by the value $\mu_{0}=-0.025$ for $N_{r}=10^{5}$ while $\mu_{0}=-0.028$ for $N_{r}=10^{6}$.
@@ IV Results for competition of three groups of entries
From $\eta_{C}(i)$ we compute for $C\in\{R,G,B\}$ its global network average (over non-white nodes) which is noted $\eta_{0,C}$ and we characterize the node preference for color $C$ by the difference $\Delta\eta_{C}(i)=\eta_{C}(i)-\eta_{0,C}$ which represents the color preference of node $i$ in comparison to the global network color preference (both for color $C$).
@@
Concerning the statistical error of $\eta_{C}(i)$ or $\Delta\eta_{C}(i)$, we mention that the theoretical error can be obtained in a similar way as for the case of two group competitions. Now, we use that $\eta_{C}(i)$ is the average (over the $N_{r}$ random pathway realizations) of $\sigma_{i}(C)$ a quantity which has values 0 or 1 such that $\sigma_{i}(C)^{2}=\sigma_{i}(C)$. This allows to compute the variance from its average and gives the theoretical error of $\eta_{C}(i)$ as $\sqrt{\eta_{C}(i)[1-\eta_{C}(i)]/(N_{r}-1)}\approx 1/(2\sqrt{N_{r}})\approx 0.0015$ for $N_{r}=10^{5}$ and $\eta_{C}(i)\approx 0.5$ (value of maximal theoretical error) which is similar to the two color case. (The factor 1/21/21 / 2 is a trivial effect of the formula $\eta_{C}(i)=(1\pm\mu)/2$ for the two color case.) We have also verified by the method of sample averages that the error of $\Delta\eta_{i}(C)$ is typically reduced by the same factor 232-32 - 3 as for the two color case.

### v3

@@ I Introduction
The process of opinion formation in human society gains higher and higher importance with the development of social networks which start to produce an important impact on political views and elections (see e.g. [1, 2]). The statistical properties of such social networks typically have a scale-free structure as reviewed in [3, 4]. Various voter models have been proposed and studied by different groups with a development of physical concepts and their applications to sociophysics [3, 5, 6, 7, 8, 9, 10, 11].
Recently, we proposed the Ising Network Opinion Formation (INOF) model and analyzed its applications to Wikipedia networks for 6 language editions of 2017 [12]. This model allows to determine opinion polarization for all Wikipedia articles (or nodes) induced by two groups of nodes with fixed opposite opinions (red or blue, spin up or down). In this INOF model the initial two groups of one or a few nodes have fixed opposite opinions represented by spin up (red color) or spin down (blue color). All other nodes have initially an undecided opinion (spin zero or white color). The formation of a steady-state opinion of each node emerges as a result of an asynchronous Monte Carlo process in which an opinion of a given node $i$ is determined by a majority vote of his friends presented by spins up or down or zero from all network nodes $j$ that have links to node $i$. Such spin flips, induced by local majority votes, are done for all nodes without repetitions in random order over all $N$ nodes. These procedure is repeated up to convergence to a steady-state for a sufficiently long time $\tau$ and corresponds to a particular random pathway realization for the order of spins to be flipped. Finally, an average over a high number of pathways realizations is done to obtain averages and distributions of the opinions for nodes or the whole network (see next section for more technical details). A somewhat similar procedure is used in the studies of problems of associative memory (see e.g. [13, 14]) even if there are significant differences from the INOF model due to the absence of certain fixed nodes and other initially white nodes and the use of positive/negative transition elements between nodes while all of them are positive for the INOF case considered here.
The mathematical and statistical properties of certain models of Ising spins om complex networks have been analyzed in [15, 16]. However, these works studied very different aspects of such networks compared to those discussed in our work.
@@
In this work, we extend the studies of the INOF model [12] to more recent Wikipedia networks collected either in October 2024 or collected at 20 March 2025 and to new specific initial groups with fixed opinions. For example, we analyze the competition between Apple Inc. vs. Microsoft, Donald Trump vs. Vladimir Putin and others. Furthermore, the competition of three entries is also considered for several cases. Thus we analyze the competition between three groups of Donald Trump, Vladimir Putin, Xi Jinping, and compare the results with the case of competition between 3 related countries USA, Russia, China. We also show that the INOF approach allows to analyze the interaction and influence of such social concepts as Liberalism, Communism, Nationalism. In this work we consider two types of vote contributions (determined by two representations of $V_{ij}$ in (1), see below) while only one of them was considered in [12]. And finally, we also examine the effects of fluctuations appearing as a result of a certain effective temperature in the voting process.
@@
We note that the preliminary results of this research have been presented at the contributed talk of authors at Wiki Workshop 2025 on May 22 (the abstract of 3 pages is available at [23]).
@@ II Model description and data sets
In this work, we use mostly three very recent Wikipedia editions (English EN, Russian RU, Chinese ZH) collected at 20 March 2025 with the number of network nodes/articles N=6969712, 2035086, 1468935N=6969712,\,2035086,\,1468935 and the number of links $N_{\ell}=190031938,\,44188839,\,21160179$ for EN, RU, ZH respectively. For certain cases, we also use the English (EN) and French (FR) Wikipedia network collected at 1 October 2024 and already used previously (see [23] and Refs. therein) with $N=6891535,2638634$ and $N_{\ell}=185658675,76118849$ for EN (2024), FR (2024) respectively.
Wikipedia articles correspond to network nodes and citation from a given article $j$ to another article $i$ correspond to a directed link with the adjacency matrix element $A_{ij}=1$ (and $A_{ij}=0$ in absence of a link from $j$ to $i$); multiple citations from $j$ to $i$ are considered as only one link. The above definition of link directions of $A_{ij}$ corresponds to those used in our previous works, see e.g. [23] and Refs. therein, and we keep it here. Then the matrix of Markov transitions is defined by $S_{ij}=A_{ij}/k_{j}$ where $k_{j}=\sum_{i}A_{ij}$ is a number of out-going links from node $j$ to any other node $i$ (such that $A_{ij}\neq 0$); for the case of dangling nodes without out-going links (i.e. with $k_{j}=0$), we simply define $S_{ij}=1/N$ implying the usual column sum normalization $\sum_{i}S_{ij}=1$ for all $j$. For later use, we also introduce the modified matrix $\tilde{S}_{ij}$ which is identical to $S_{ij}$ for $k_{j}>0$ and with $\tilde{S}_{ij}=0$ for dangling nodes.
Usually, in other typical types of network studies (see e.g. [23]) one introduces the Google matrix of the network defined as $G_{ij}=\alpha S_{ij}+(1-\alpha)/N$ where alpha is the damping factor with the standard value $\alpha=0.85$ [24, 25]. Here the network nodes can be characterized by the PageRank vector which is the eigenvector of the Google matrix $G$ [24, 25] with the highest eigenvalue $\lambda=1$, i.e. $GP=\lambda P=P$, and the damping factor $\alpha<1$ ensures that this vector is unique and can be computed efficiently. Its components $P(i)$ are positive and normalized to unity (${\sum_{i=1}}^{N}P(i)=1$). The network nodes $i$ can be ordered by monotonically decreasing probabilities $P(i)$ which provides the PageRank index $K$ with with highest probability at $K=1$ and smallest at $K=N$. Some results for the PageRank vector and its index for recent Wikipedia editions of 2024 can be found at [23] and Refs. therein. However, in this work, we do not use the Google matrix neither the PageRank and focus mostly on the modified matrix $\tilde{S}$ and also the adjacency matrix $A_{ij}$ to define an asynchronous Monte Carlo process.
As in [12] a few selected nodes (wiki-articles) have assigned fixed spin values $\sigma_{l}=-1$ blue e.g. for Microsoft and $\sigma_{k}=1$ red for Apple Inc.. These specific spin nodes always keep their polarization. All other nodes $i$ are initially assigned with a white color (or spin $\sigma_{i}=0$) and have no definite initial opinion. However, once they acquire a different color red or blue (spin value $\sigma_{i}=\pm 1$) during the asynchronous Monte Carlo process they can flip only between $\pm 1$ and cannot change back to the white opinion.
@@
In Eq. (1) $\sigma_{j}=1$ if the spin of node $j$ is oriented up (red color), or $\sigma_{j}=-1$ if it is oriented down (blue color), or $\sigma_{j}=0$ if the node $j$ has no opinion (if it has still its initial white value). After the computation of $Z_{i}$, the spin $\sigma_{i}$ of node $i$, is updated: it becomes $\sigma_{i}=1$ if $Z_{i}>0$, $\sigma_{i}=-1$ if $Z_{i}<0$, and remains unchanged if $Z_{i}=0$. This operation is repeated for all non-fixed nodes $i^{\prime}$ following a predetermined random order (shuffle) such that there is no repetition at this level and each spin is updated only once. Note that due to the possibility of $Z_{i}=0$ it is possible that a node $i$ keeps his initial white value $\sigma_{i}=0$. After the update of $\sigma_{i}$, the modified value of $\sigma_{i}$ is used for the computation of $Z_{i^{\prime}}$ of subsequent values $\sigma_{i^{\prime}}$.
@@
The voting process for the case of a competition between three groups of entries is an extension of this procedure and its details are explained later.
@@ III.1 Comparison of OPA and OPS
Figure 1: Probability density $p(f_{r})$ of fraction of red nodes $f_{r}$ for SCCI for EN Wiki2025 versus $f_{r}$ using the matrix AA (OPA case, top panel) or the matrix $\tilde{S}$ (OPS case, bottom panel). The value of $f_{r}$ for a given pathway realization gives the fraction of red outcome of network nodes for Socialism/Communism (fraction computed with respect to all non-white network nodes). The histograms show the distribution of $f_{r}$ with respect to all $N_{r}$ different random pathway realizations and all nodes. The normalization is fixed by $\int_{0}^{1}p(f_{r})\,df_{r}=1$ with bin width $0.005$ (i.e. 200 bins in the full interval for $f_{r}\in[0,1]$). The red (green) curves correspond to $N_{r}=100000$ ($N_{r}=10000$). Note that the distribution $\tilde{p}(f_{b})$ for blue outcome for Capitalism/Imperialism (not shown in the figure) can be obtained by symmetry $\tilde{p}(f_{b})=p(1-f_{b})$ since $f_{r}+f_{b}=1$.
@@
Figure 2: Probability density $p(\mu)$ of polarization $\mu$ for SCCI for EN Wiki2025 versus $\Delta\mu=\mu-\mu_{0}$ using the matrix AA (OPA case, top panel) or the matrix $\tilde{S}$ (OPS case, bottom panel). Here a red (blue) outcome with $\mu_{i}\approx 1$ ($\mu_{i}\approx-1$) corresponds to Socialism/Communism (Capitalism/Imperialism). The histograms show the distribution of $\mu_{i}$ for all non-white nodes $i$ in the network with with $n_{r}(i)+n_{b}(i)\approx N_{r}$ and white nodes with $n_{r}(i)+n_{b}(i)\ll N$ are not taken into account. The normalization is fixed by $\int_{-1}^{1}p(\mu)\,d\mu=1$ with bin width $0.01$ (i.e. 200 bins in the full interval for $\mu_{i}\in[-1,1]$). Values of $p(\mu)$ outside the shown interval [0.2,0.2][-0.2,0.2] are zero on graphical precision. The red (green) curve corresponds to $N_{r}=100000$ ($N_{r}=10000$). The value $\mu_{0}$ is the average of $\mu_{i}$ with respect to all (non-white) nodes $i$ (computed for $N_{r}=100000$) and has the values $\mu_{0}=-0.02333$ for OPA and $\mu_{0}=-0.10886$ for OPS.
@@
It is also possible to compute the statistical error quite accurately by a more direct numerical computation. For this we divide the full data of global $N_{r}=10^{5}$ pathway realizations into 100 samples with reduced $N_{r}=10^{3}$ and compute for each sample partial averages $\mu_{i}$, $\mu_{0}$ and $\Delta\mu_{i}$ over the reduced value of $N_{r}=10^{3}$. This provides in particular partial averages $\mu_{0}$ with quite significant fluctuations between the samples. Using the 100 partial average values $\Delta\mu_{i}$ for each sample it is straightforward to determine the statistical error of $\Delta\mu_{i}$ as $\approx\sqrt{\langle\Delta\mu_{i}^{2}\rangle-\langle\Delta\mu_{i}\rangle)^{2}/(100-1)}$ where $\langle\ldots\rangle$ is the simple average over the 100 samples (of the partial sample averages). It turns out that typical error values obtained in this way (at global $N_{r}=10^{5}$ and for SCCI of EN Wiki2025 and a few other cases) are closer to 0.00080.00150.0008-0.0015 which is about 232-3 times smaller than the theoretical error $\approx 0.003$. This reduction is apparently due to rather strong correlations between $\mu_{i}$ and $\mu_{0}$ which are likely to have statistical fluctuations in the same direction (partial sample averages of both are likely to be “large” or “small” at the same time). These correlations are also visible in the two peak distribution of Figure 1 showing that in one given pathway realization it is rather likely to have either $f_{r}\approx 1$ (or $f_{r}\approx 0$) corresponding to a majority of network nodes $i$ with either $\sigma_{i}=1$ (or $\sigma_{i}=-1$ respectively). We note that the numerical statistical error of $\mu_{0}$, which is obtained as byproduct of the above procedure, is $\approx 0.002-0.003$ which is closer to the theoretical error.
@@
We also consider the INOF competition between Emmanuel Macron ($\mu=1$) vs. Marine Le Pen ($\mu=-1$) with $\mu_{0}=-0.028$ from FR Wiki2024. These corresponds to rather close vote fractions with $f_{r}=(1+\mu_{0})/2=0.4860$ for Macron and $f_{b}=(1-\mu_{0})/2=0.5140$ for Le Pen (number of pathway realizations $N_{r}=10^{6}$). The polarization opinions of 14 French political figures and top 10 French richest persons from the Forbes list 2015-2024 are presented in Appendix Table A1. The results for these 24 persons show that only two of them have a polarization in favor of Emmanuel Macron and other 22 in favor of Marine Le Pen. At the same time the values of $\Delta\mu$ are mainly located at relatively small values $|\Delta\mu|\sim 0.01$ being approximately by a factor 10 smaller of values of for the contest of Trump-Putin shown in Table 1. Due to these smaller $\Delta\mu$ values we also compute an additional data set with $N_{r}=10^{6}$ to reduce the statistical errors (note that the FR Wiki2024 network is roughly 2 times smaller in number of nodes and links as compared to those of EN Wiki2025 thus reducing the numerical effort). Table A1 shows two columns of $\Delta\mu$ computed for $N_{r}=10^{5}$ and $N_{r}=10^{6}$ (the smaller data set is statistically independent and not included in the larger date set). The values are rather close and the small differences indicate the size of the typical fluctuations at $N_{r}=10^{5}$.
The theoretical error of the $\mu_{i}$ values is $(1-\mu_{i}^{2})/\sqrt{N_{r}}\approx 0.003$ (for $N=10^{5}$) or $\approx 0.001$ (for $N=10^{6}$). However, the more precise error estimation for $\Delta\mu_{i}=\mu_{i}-\mu_{0}$ using a subdivision of the data in 100 samples (described above after the discussion for Figure 2) provides typical statistical errors of $\Delta\mu_{i}$ for the entries in Table A1 which are roughly 3 times smaller than the theoretical error, i.e. $\approx 0.001$ (for $N=10^{5}$) or $\approx 0.0003$ (for $N_{r}=10^{6}$) which is due to rather strong correlations between $\mu_{i}$ and $\mu_{0}$. Therefore, despite the typical small values of $\Delta\mu_{i}\approx 0.01$ in Table A1 their relative error is mostly only 3% (for the $N_{r}=10^{6}$ data). Note that the error $0.0023$ of $\mu_{0}$ itself is closer to the theoretical error $0.003$ which is confirmed by the value $\mu_{0}=-0.025$ for $N_{r}=10^{5}$ while $\mu_{0}=-0.028$ for $N_{r}=10^{6}$.
@@ IV Results for competition of three groups of entries
As in the case of competition of 3 currencies in [18] the competition of 3 entries allows to highlight more complex interactions between 3 selected entries compared to the case of only 2 entries.
@@
From $\eta_{C}(i)$ we compute for $C\in\{R,G,B\}$ its global network average (over non-white nodes) which is defined as $\eta_{0,C}$ and we characterize the node preference for color $C$ by the difference $\Delta\eta_{C}(i)=\eta_{C}(i)-\eta_{0,C}$ which represents the color preference of node $i$ in comparison to the global network color preference (both for color $C$).
@@
Concerning the statistical error of $\eta_{C}(i)$ or $\Delta\eta_{C}(i)$, we mention that the theoretical error can be obtained in a similar way as for the case of two group competitions. Now, we use that $\eta_{C}(i)$ is the average (over the $N_{r}$ random pathway realizations) of $\sigma_{i}(C)$ a quantity which has values 0 or 1 such that $\sigma_{i}(C)^{2}=\sigma_{i}(C)$. This allows to compute the variance from its average and gives the theoretical error of $\eta_{C}(i)$ as $\sqrt{\eta_{C}(i)[1-\eta_{C}(i)]/(N_{r}-1)}\approx 1/(2\sqrt{N_{r}})\approx 0.0015$ for $N_{r}=10^{5}$ and $\eta_{C}(i)\approx 0.5$ (value of maximal theoretical error) which is similar to the two color case. (The factor $1/2$ is a trivial effect of the formula $\eta_{C}(i)=(1\pm\mu)/2$ for the two color case.) We have also verified by the method of sample averages that the error of $\Delta\eta_{i}(C)$ is typically reduced by the same factor 232-3 as for the two color case.